Sheri's age in 20 years will be the same as Terry's age now. Ten years from now, Terry's age will will be twice as Sheri's. What are their present ages?
Answer by stanbon(75887)

Sheri's age in 20 years will be the same as Terry's age now. Ten years from now, Terry's age will will be twice as Sheri's. What are their present ages?
--------------------
NOW DATA:
Let Sheri's age be x yrs ; Let Terry's age be y yrs old.
------------------
20 yrs from NOW DATA:
Sheri will be x+20 ; Terry will be y+20
----------------
10 yrs from NOW DATA:
Sheri will be x+10, Terry will be y+10
-----------------------
EQUATIONS:
x+20 = y
y+10=2(x+10)
---------------
Rearrange and simplify to get:
1st: x - y = -20
2nd: 2x - y = -10
---------------------
Subtract 1st from 2nd to get:
x = 10 (Sheri's age NOW)
Substitute into y = x+20 to solve for y:
y = 10+20 = 30 (Terri's age NOW)
----------------
Cheers,
Stan H.